OpenStudy (anonymous):

What is the tangent line of y = sqrt(x-1)? Please show steps thanks

OpenStudy (campbell_st):

you can only get the equation of the gradient of the tangent from differentiating. You will need a point or ordered pair to find the equation of the tangent. rewrite the equation in index form \[y = (x + 1)^{\frac{1}{2}}\] the differentiating using the chain rule \[\frac{dy}{dx} = \frac{1}{2} ( x + 1)^{-\frac{1}{2}}\] or \[\frac{dy}{dx} = \frac{1}{2\sqrt{x + 1}}\] but this is only the equation of the gradient of the tangent.... you still need an ordered pair or x value for a specific tangent

OpenStudy (anonymous):

ok for the ordered pair (2,1)

OpenStudy (campbell_st):

ok to find the gradient of the tangent at x = 2 just substitute x =2 \[m = \frac{1}{2\sqrt{2 + 1}}... m = \frac{1}{2\sqrt{3}}\] then using the point slope formula \[y - 1 = \frac{1}{2\sqrt{3}}( x - 2)\] will be the equation of the tangent at x = 2

OpenStudy (campbell_st):

oops you'll need to simplify things a bit

OpenStudy (anonymous):

So I can just do the power rule to find the f^1(x) and that will give me the slope?

OpenStudy (campbell_st):

it gives the equation of the slope... you still need an x value to get the slope at a point...

OpenStudy (anonymous):

ok and once you have the slope how do you put it into the linear form? can you do y = mx+b?

OpenStudy (campbell_st):

yes you can... once you know m.... put it into slope intercept form.. substitute the ordered pair to find the intercept.

OpenStudy (anonymous):

so is the slope at (2,1) 2?

OpenStudy (campbell_st):

no its \[m = \frac{1}{2\sqrt{3}}\] this comes from substituting x = 2 into the 1st derivative

OpenStudy (anonymous):

why is it 1/2(x+1)^(-1/2) instead of 1/2(x-1)^(-1/2) if the original function was f(x) = (x-1)^(-1/2)?

OpenStudy (campbell_st):

the original function you posted was \[y = \sqrt{x + 1}... or... y = (x+1)^{\frac{1}{2}}\] sorry but I have to go... good luck with it

OpenStudy (anonymous):

no I posted y = sqrt(x-1) but okay